What are all the solutions to -10/(x-9) = x/2
Answer by The One(11)

-10/(x-9)=x/2

Since there is one rational expression a mathematical symbol, or combination of symbols, representing a value, or relation. Example: 2+2=4 on each side of the equation a mathematical statement that says two expressions have the same value; any number sentence with an = . , this can be solved as a ratio a pair of numbers that compares different types of units. . For example, (AĤB=C/D) is equivalent two or more expressions that have the same value. To A*D=B*C.
-10 • 2=x • (x-9)
Since x is on the right-hand side of the equation a mathematical statement that says two expressions have the same value; any number sentence with an = . , switch the sides so it is on the left-hand side of the equation a mathematical statement that says two expressions have the same value; any number sentence with an = . .
x • (x-9)=-10 • 2
Multiply to compute a product; to perform a multiplication. x by (x-9) to get (x-9).
x(x-9)=-10 • 2
Multiply to compute a product; to perform a multiplication. -10 by 2 to get -20.
x(x-9)=-20
Multiply to compute a product; to perform a multiplication. x by each term any expression written as a product or quotient. Example: 2xy , 4m^2, or -3x^2y^3z inside the parentheses marks of inclusion (symbols: ( and ) ). A parenthesis is the plural form of parenthesis. .
x^2-9x=-20
To set a well-defined group of objects. the left-hand side of the equation mathematical statement that says two expressions have the same value; any number sentence with an = . equal to 0, move all the expressions a mathematical symbol, or combination of symbols, representing a value, or relation. Example: 2+2=4 to the left-hand side.
x^2-9x+20=0
In this problem -4*-5=20 and -4-5=-9, so insert -4 as the right hand term any expression written as a product or quotient. Example: 2xy , 4m^2, or -3x^2y^3z of one factor one of two or more expressions that are multiplied together to get a product. and -5 as the right-hand term any expression written as a product or quotient. Example: 2xy , 4m^2, or
-3x^2y^3z of the other factor one of two or more expressions that are multiplied together to get a product. .
(x-4)(x-5)=0
Set a well-defined group of objects. each of the factors one of two or more expressions that are multiplied together to get a product. of the left-hand side of the equation. A mathematical statement that says two expressions have the same value; any number sentence with an = . equal to 0.
x-4=0
x-5=0

Since -4 does not contain the variable. A letter used to represent a number value in an expression or an equation. to solve for, move it to the right-hand side of the equation. A mathematical statement that says two expressions have the same value; any number sentence with an = . by adding 4 to both sides.
x=4
x-5=0
Set a well-defined group of objects. each of the factors. One of two or more expressions that are multiplied together to get a product. of the left-hand side of the equation. A mathematical statement that says two expressions have the same value; any number sentence with an = . equal to 0.
x=4
x-5=0
Since -5 does not contain the variable. A letter used to represent a number value in an expression or an equation. to solve for, move it to the right-hand side of the equation. A mathematical statement that says two expressions have the same value; any number sentence with an = . by adding 5 to both sides.
x=4
x=5
The complete solution. The value of a variable that makes an equation true. is the set a well-defined group of objects. of the individual solutions. The value of a variable that makes an equation true. .
x=4,5